OpenStudy (anonymous):

Brandon wants to retire in 42 years. He makes annual contributions to a Traditional IRA in the amount of $2,200. He is taxed at 33% today, but anticipates that will change to 15% in his retirement. If the average annual rate of return on the account is 4.9%, what is the effective value of Brandon’s IRA account when he retires? @ganeshie8

ganeshie8 (ganeshie8):

annuity setup

ganeshie8 (ganeshie8):

do u have the formula for future value ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

one sec, let me pull up my notes

ganeshie8 (ganeshie8):

\(\large FV = C\bullet \frac{(1+i)^{nt}-1}{i}\)

ganeshie8 (ganeshie8):

plug the numbers and find out the future value, we can subtract the tax to get the effective value

ganeshie8 (ganeshie8):

C = periodic cash payments = 2200 i = r/n = 0.049/1 = 0.049 n = 1 t = 42

ganeshie8 (ganeshie8):

plugin

ganeshie8 (ganeshie8):

once u have the FV, simply subtract the tax of 15% of the FV

ganeshie8 (ganeshie8):

im getting late for dinner.... brb